OpenStudy (anonymous):

solve the complex number system (x^3)-125=0 in a+bi form

OpenStudy (triciaal):

do you recognize the difference of cubes?

OpenStudy (triciaal):

a^3 - b^3 = (a - b)( a^2 + ab + b^2) you can factor the quadratic ( a + b)^2 - ab =0 (a + b)^2= ab a + b = =/-sqrt ab for this problem a = x and b = 5

OpenStudy (campbell_st):

well there is no complex solution if you rewrite it you get \[x^3 = 125\] so x = 5

OpenStudy (campbell_st):

all you need to do is take the cube root of both sides

OpenStudy (triciaal):

@campbell_st in a+bi form factoring special products ----difference of cubes I did something like this recently it has to to with the + and - root

OpenStudy (triciaal):

@satellite73 will you please help

OpenStudy (campbell_st):

ok... I'd say x = 5 + 0i